Be sure to answer all parts. The reaction 2A ri
Answer #1

Part A

second order equation

1/[At] = kt + [A0]

where [A0] = intial concentration = 0.00779 M

[At] = concentrationat certain time = 0.0018

k = rate constant = 32.1 M-1 s-1

put all the values in the above equation

1/0.0018 = 32.1 x t + 1/0.00779

555.56 = 32.1 x t + 128.37

2.1 x t = 555.56 - 128.37

2.1 x t = 427.19

t = 427.19 / 2.1

= 203.424 sec

Part B

t1/2 = 1/k [A0]

= 1/ 32.1 M-1 s-1x 0.00779M

= 1/0.25

= 4 s

part C

t1/2 = 1/k [A0]

= 1/ 32.1 M-1 s-1x 0.00269M

= 1/0.0863

= 11.58 s